well i`ve been trying this problem for hours "A field has a perimeter of 240 yards. The length is 20 yards less than the width. Find the length and width"?
I cant find the answer.

Answer by checkley71(8403)

1240=2*W+2(W-20)
240=2W+2W-40
4W=240+40
4W=280
W=280/4
W=70 YARDS WIDE THEN THE LENGTH IS
70-20=50 YARDS FOR THE LENGTH
PROOF
240=2*70+2*50
240=140+100
240=240